An Estes toy rocket accelerates in a straight line direction from the ground at 27.0 m/s2, and at an angle of 53.0° above the horizontal while the engine is burning fuel. The rocket fuel burns out after 3.7 s and the rocket takes a projectile path to the ground. Find its horizontal distance from the launch point to where it hits the ground. in M
Answer #1

for 3.78 Yo-Yi t Vigt 5 ( 87 sin 53°) (3.732 h (0) + 2 hi 147.6 m 2 1 and Sx= xi-o (27 cas 53) (9.12) 2 9x1 x 111.22 m and 64.98² 79.8t - 14706 = 0 t: 17.968 After this time, the it will hit ground. {as 9x=0} So sxe - Vix) X- 111-22 - (60-12) (17.36